OpenStudy (anonymous):

Find the term x^3 in the expansion of (2/3 x - 3) ^8 @strawberryswing @rvc @ganeshie8 @mathstudent55

OpenStudy (anonymous):

U found me in math chat aye

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

yes, can you please help :) @strawberryswing @Nnesha

OpenStudy (anonymous):

or do you know anyone that can help me? @strawberryswing @Nnesha

OpenStudy (anonymous):

@arabbride

OpenStudy (anonymous):

Yeah idk

OpenStudy (anonymous):

Ask hmmm @Jhannybean

OpenStudy (anonymous):

@SyedMohammed98

OpenStudy (jhannybean):

Oh, nevermind. This is a Binomial Expansion. Sorry, Im not too familiar with these. @mathstudent55 @mathmath333

OpenStudy (jhannybean):

Hm... I could try,

OpenStudy (anonymous):

I'll appriaciate :) @Jhannybean however im really confused

OpenStudy (mathmath333):

first find the terms expansion of \((a-b)^8=a^8-a^7b+a^6b^2-a^5b^3+a^4b^4-a^3b^5+a^2b^6-ab^7+b^8\) and see the the coefficients by the pascal triangle 1 1 2 1 1 3 3 1 1 4 6 4 1 1 5 10 10 5 1 1 6 15 20 15 6 1 1 7 21 35 35 21 7 1 1 8 28 56 70 56 28 8 1

OpenStudy (anonymous):

what do i plug in a and b? @mathmath333

OpenStudy (mathmath333):

here \(a=\dfrac{2x}{3}~~and~~b=3\)

OpenStudy (jhannybean):

your function.

OpenStudy (mathmath333):

first find where is the expansion is \(\huge a^3\)

OpenStudy (mathmath333):

* in the expansion of \((a-b)^8\)

OpenStudy (anonymous):

so for (a-b)^8 i plug in ( ( 2x/3) - (3) ^8 ?

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@Jhannybean

OpenStudy (mathmath333):

no need to do that

OpenStudy (mathmath333):

u want to find the term of \(x^3\) right ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

so we can see that in the expansion of \((a-b)^8\) there is term \(-a^3b^5\) which is the sixth term and in the pascal triangle in the the eight row of \(6^{th}\) term is \(56\) so ur required term is \(-56a^3b^5\) now plug \(a=\dfrac{2x}{3} \) and \(b=3\)

OpenStudy (anonymous):

-56(2x/3)(3^5) ? @mathmath333

OpenStudy (mathmath333):

yes simplify it

OpenStudy (mathmath333):

actually it is\(\huge -56(\dfrac{2x}{3})^3(3^5)\)

OpenStudy (mathmath333):

u forgot the power of \(\dfrac{2x}{3}\)

OpenStudy (mathmath333):

there is also a much shorter way for this

OpenStudy (anonymous):

i honestly dont know what to do from here -56 (2x/3) ^3 (243)

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

u can \((\dfrac{2x}{3})^3\\ =\dfrac{2^3x^3}{3^3} \)

OpenStudy (mathmath333):

this will be your final result https://www.wolframalpha.com/input/?i=-56%28%5Cdfrac%7B2x%7D%7B3%7D%29%5E3%283%5E5%29

OpenStudy (anonymous):

thank you so much @mathmath333 and everyone else who tried to help :) i dont want to be abusing of your assistnance but can you helo me in this problem too let S be the total area of the two segments shaded in the diagram below  (c) show that S=2 (pie - 2 2 sin 0 ) (d) find the value of 0 when s is a local mininum,justifying that it is a minimum (e) find the value of 0 for which S has its greates value.